OpenStudy (anonymous):

Which statement is true for the equation 2x − 2x − 7 = −7?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is an equation that is true

OpenStudy (anonymous):

hi!

OpenStudy (misty1212):

since \(2x-2x=0\) you really have \(-7=-7\) so it is an identity as well

OpenStudy (anonymous):

? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (anonymous):

those are the possible answers

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@strawberryswing

OpenStudy (anonymous):

Do you know where to begin?

OpenStudy (anonymous):

x - 7= -7 x=0

OpenStudy (anonymous):

right??

OpenStudy (anonymous):

Since the equation is 2x-2x-7=-7, you would take out the 2x-2x, since you would get 0. and you are left with -7=-7

OpenStudy (anonymous):

So you would only have one solution. Since -7=-7..

OpenStudy (anonymous):

its correct thank you

OpenStudy (anonymous):

You're welcome.